# Polycyclic codes over serial rings and their annihilator CSS construction

### v1

@@ 7 Annihilator CSS construction
The CSS construction is a well-known method to construct quantum codes from linear codes over fields or rings, see for example [17, 29, 33]. Depending on the duality defined for the code, this technique is referred to as either the Euclidean CSS construction or the Hermitian CSS construction. As we noticed above, polycyclic codes are also equipped with the annihilator dual. In this section, we establish annihilator CSS construction, a novel approach for constructing quantum codes from annihilator dual-preserving polycyclic codes over finite chain rings.
Recall that since $R$ is a finite commutative chain ring with identity it is a Frobenius ring, thus it has generating characters. For the character $\chi$ in $\operatorname{Hom}(R;\mathbb{C}^{*})$, there exists a unique function $\psi$ in $\operatorname{Hom}(R;\mathbb{Q}/\mathbb{Z})$ such that $\chi(x)=\exp(2\pi i\psi(x))$. Let $\langle\,,\,\rangle_{\chi}:{R}^{2n}\times{R}^{2n}\rightarrow\mathbb{Q}/\mathbb{Z}$ and $\langle\,,\,\rangle_{s}:{R}^{2n}\times{R}^{2n}\rightarrow{R}$ be two forms defined as follows:
$\langle(a|b),(a^{\prime}|b^{\prime})\rangle_{\chi}=\psi(b\cdot a^{\prime}-b^{\prime}\cdot a)\quad\text{and}\quad\langle(a|b),(a^{\prime}|b^{\prime})\rangle_{s}=b\cdot a^{\prime}-b^{\prime}\cdot a.$
We denote by $S^{\perp_{\chi}}$ and $S^{\perp_{s}}$ as the sets:
$S^{\perp_{\chi}}=\{u\in{R}^{2n}\,|\,\langle(s,u)\rangle_{\chi}=0,\,\forall s\in S\}\quad\text{and}\quad S^{\perp_{s}}=\{u\in{R}^{2n}\,|\,\langle(s,u)\rangle_{s}=0,\,\forall s\in S\}.$
Lemma 3 ([29] Lemma 6).
Let $R$ be a finite commutative chain ring and $S$ be a subgroup of $(R^{2n},+)$. Then $|S|\,|S^{\perp_{\chi}}|=|R^{2n}|$.
Lemma 4 ([29] Lemma 7).
Let $R$ be a finite commutative chain ring and $S$ and $T$ be $R$-submodules of $R^{2n}$. Then $S\perp_{\chi}T$ if and only if $S\perp_{s}T$.
A quantum code $Q$ of length $n$ and size $K$ is a $K$-dimensional subspace of the $\mathfrak{q}^{n}$-dimensional Hilbert space $\mathbb{C}^{\mathfrak{q}^{n}}=\mathbb{C}^{\mathfrak{q}}\otimes\cdots\otimes\mathbb{C}^{\mathfrak{q}}$. A code with a minimum distance $d$ can detect errors that affect up to $d-1$ subsystems non-trivially. Moreover, it can correct errors affecting up to $1/2(d-1)$ of the subsystems. Quantum codes are often denoted with $((n,K,d))_{\mathfrak{q}}$, where $n$, $K$, and $d$ are as defined above. Let $R$ have $\mathfrak{q}$ elements and $B$ be an orthonormal basis of $\mathbb{C}^{\mathfrak{q}}$ concerning the usual Hermitian inner product. The elements of the basis $B$ are labeled by ring elements, such that $B=\{|x\rangle\,|\,x\in R\}$, where ||\,$\rangle$| ⟩ is Dirac ket notation. For each $a\in R$ and $b\in R$, the operator $X(a):\mathbb{C}^{\mathfrak{q}}\rightarrow\mathbb{C}^{\mathfrak{q}}$ is defined by $X(a)|x\rangle=|x+a\rangle$, and the operator $Z(b):\mathbb{C}^{\mathfrak{q}}\rightarrow\mathbb{C}^{\mathfrak{q}}$ is defined as $Z(b)|x\rangle=\chi(bx)|x\rangle$, for all $x$ in $R$. These operators can be extended for the space $\mathbb{C}^{\mathfrak{q}^{n}}$ as $X(a)=X(a_{1})\otimes X(a_{2})\otimes\cdots\otimes X(a_{n})$ and $Z(b)=Z(b_{1})\otimes Z(b_{2})\otimes\cdots\otimes Z(b_{n})$, where $a=(a_{1},\ldots,a_{n})\in{R}^{n}$ and $b=(b_{1},\ldots,b_{n})\in{R}^{n}$. Let $G_{n}$ be a group consisting of all $\chi(c)X(a)Z(b)$, where $a,b\in R^{n}$ and $c\in R$. For any subgroup $S$ of $G_{n}$, $\text{Fix}(S)$ is defined as
$\text{Fix}(S)=\{v\in\mathbb{C}^{\mathfrak{q}^{n}}\mid Ev=v\text{ for all }E\in S\}.$
The quantum code $Q$ is called stabilizer code if and only if there exists a subgroup $S$ of $G_{n}$ such that $\text{Fix}(S)=Q$.
The symplectic weight of $(a|b)\in R^{2n}$, denoted by $\operatorname{swt}((a|b))$, is defined as the number of indices $i$, where $a_{i}\neq 0$ or $b_{i}\neq 0$. In other words, $\operatorname{swt}((a|b))=\left|\{k\,|\,(a_{k},b_{k})\neq(0,0)\}\right|$. It is well-known that
$\frac{1}{2}\operatorname{wt}((a|b))\leqslant\operatorname{swt}((a|b))\leqslant\operatorname{wt}((a|b)),$ (21)
where $\operatorname{wt}((a|b))$ denotes the Hamming weight of the vector $(a|b)$, see [35].
Theorem 13 ([29] Theorem 8).
An $((n,K,d))_{\mathfrak{q}}$ stabilizer code exists if and only if there exists an additive code $C\subseteq R^{2n}$ of size $|C|=|R|^{n}/K$ such that $C\subseteq C^{\perp_{\chi}}$ and $\operatorname{swt}(C^{\perp_{\chi}}\backslash C)=d$ if $K>1$, and $\operatorname{swt}(C^{\perp_{\chi}}\backslash\{0\})=d$ if $K=1$.
A linear code over the ring $R$ of length $n$, rank $k$, and minimum Hamming distance $d$ is denoted by the parameters $[n,k,d]_{R}$. Now, we are ready to present an annihilator counterpart for the Euclidean and Hermitian CSS construction, as presented in [17, 29, 33].
@@
Suppose that $R$ is a chain ring and $f(x)$ defined as in Equation (3) is a polynomial in $R[x]$. Let $f_{0}\in R$ be an invertible element, $f(x)=g_{1}(x)h_{1}(x)=g_{2}(x)h_{2}(x)$, $C_{1}=\langle g_{1}(x)\rangle$ and $C_{2}=\langle g_{2}(x)\rangle$ be two $f(x)$-polycyclic codes over the ring $R$ with parameters $[n,k_{1},d_{1}]_{R}$ and $[n,k_{2},d_{2}]_{R}$, respectively, such that $C^{\circ}_{2}\subseteq C_{1}$. Then, there exists an $((n,k_{1}+k_{2}-n,d))_{\mathfrak{q}}$ stabilizer code with the minimum distance $d=\min\{\operatorname{wt}(c)\,|\,c\in(C_{1}\backslash C_{2}^{\circ})\cup(C_{2}\backslash C_{1}^{\circ})\}$.
@@
We define $C=C^{\circ}_{1}\times C^{\circ}_{2}\subseteq{R}^{2n}$, where $C_{1}^{\circ}$ and $C_{2}^{\circ}$ are annihilator duals of $C_{1}$ and $C_{2}$, respectively. If $(a|b)$ and $(a^{\prime}|b^{\prime})$ are in $C$, then $b\cdot a^{\prime}-b^{\prime}\cdot a=0$. Thus $C\subseteq C^{\perp_{s}}$, that is $C\perp_{s}C$. Using Lemma 4, we have $C\subseteq C^{\perp_{\chi}}$. Furthermore, if $(a|b)\in C_{2}\times C_{1}$, we have $b\cdot a^{\prime}-b^{\prime}\cdot a=0$ for any arbitrary element $(a^{\prime}|b^{\prime})$ in $C$, and hence $(a|b)\in C^{\perp_{s}}$. So, $(C_{2}\times C_{1})\perp_{s}C$. Applying Lemma 4, $(C_{2}\times C_{1})\perp_{\chi}C$, which implies $(C_{2}\times C_{1})\subseteq C^{\chi}$. On the other side, by Lemma 3 and Part (3) in Theorem 10, we have
$|C^{\perp_{\chi}}|=\frac{|R|^{2n}}{|C|}=\frac{|R|^{n}}{|C^{\circ}_{1}|}\frac{|R|^{n}}{|C^{\circ}_{2}|}=|C_{1}||C_{2}|=|C_{2}\times C_{1}|.$
Therefore, $C_{2}\times C_{1}=C^{\perp_{\chi}}$. Moreover,
$|C|=|C^{\circ}_{1}||C^{\circ}_{2}|=\frac{|R|^{n}}{|C_{1}|}\frac{|R|^{n}}{|C_{2}|}=\frac{|R|^{2n}}{|R|^{k_{1}+k_{2}}}=\frac{|R|^{n}}{|R|^{k_{1}+k_{2}-n}}.$
If we take $K=|{R}|^{k_{1}+k_{2}-n}$, then $|C|=|R|^{n}/K$. We claim that $\operatorname{swt}(C^{\perp_{\chi}}\backslash C)=d$. To prove, assume that $(a|b)\in(C^{\perp_{\chi}}\backslash C)=((C_{2}\times C_{1})\backslash(C^{\circ}_{1}\times C^{\circ}_{2}))$ is an arbitrary element. So, $(a|b)\in C_{2}\times C_{1}$ and $(a|b)\notin C^{\circ}_{1}\times C^{\circ}_{2}$. That is $a\notin C^{\circ}_{1}$ or $b\notin C^{\circ}_{2}$. Therefore, $a,b\in(C_{1}\backslash C_{2}^{\circ})\cup(C_{2}\backslash C_{1}^{\circ})$. According to the definition of $d$, $\operatorname{wt}(a)\geqslant d$ and $\operatorname{wt}(b)\geqslant d$. Applying Equation (21), we have
$\operatorname{swt}(a|b)\geqslant\frac{1}{2}\operatorname{wt}(a|b)=\frac{1}{2}(\operatorname{wt}(a)+\operatorname{wt}(b))\geqslant d.$
Thus, $\operatorname{swt}(C^{\perp_{\chi}}\backslash C)\geqslant d$. On the other side, we know there exists an element $c$ in $(C_{1}\backslash C_{2}^{\circ})\cup(C_{2}\backslash C_{1}^{\circ})$ such that $\operatorname{wt}(c)=d$. In the first case, let $c\in(C_{1}\backslash C_{2}^{\circ})$. Then we have $(0|c)\in(C^{\perp_{\chi}}\backslash C)$, and $\operatorname{swt}(0|c)=\operatorname{wt}(c)=d$. In the second case, let $c\in(C_{2}\backslash C_{1}^{\circ})$. Then we have $(c|0)\in(C^{\perp_{\chi}}\backslash C)$, and $\operatorname{swt}(c|0)=\operatorname{wt}(c)=d$. Hence in both cases, there exists an element in $(C^{\perp_{\chi}}\backslash C)$ with the symplectic weight $d$. Therefore, $\operatorname{swt}(C^{\perp_{\chi}}\backslash C)\leqslant d$. To sum up, we have found the additive code $C\subseteq R^{2n}$ of size $|C|=|R|^{n}/K$ such that $C\subseteq C^{\perp_{\chi}}$ and $\operatorname{wt}(C^{\perp_{\chi}}\backslash C)=d$ As a result, Theorem 13 implies that there exists an $((n,K,d))_{\mathfrak{q}}$ stabilizer code. ∎
@@
Suppose that $R$ is a chain ring and $f(x)$ defined as in Equation (3) is a polynomial in $R[x]$. Let $f_{0}\in R$ be an invertible element and $f(x)=g(x)h(x)$. If $C$ is an $f(x)$-polycyclic code over the chain ring $R$ with parameters $[n,k,d]_{R}$, generated by $g(x)$, and $C^{\circ}\subseteq C$, then there exists an $((n,2k-n,\geqslant d))_{\mathfrak{q}}$ stabilizer code.
Now, we are prepared to give the result that allows to construct quantum codes from $f(x)$-polycyclic codes over the serial ring $\mathscr{R}$.
Theorem 15.
Suppose that $f(x)$ defined as in Equation (3) is a polynomial in $\mathscr{R}[x]$, $f_{0}\in\mathscr{R}$ is an invertible element, and $f(x)=g(x)h(x)$. Let $\mathscr{C}$ be an annihilator dual-preserving $f(x)$-polycyclic code of length $n$ over $\mathscr{R}$, decomposed as $\mathscr{C}\cong\oplus_{i=1}^{k}\mathscr{C}_{i}$. If each $\mathscr{C}_{i}$ has parameters $[n_{i},k_{i},d_{i}]$, then there exists a stabilizer code with parameters $((\sum_{i=1}^{k}n_{i},\sum_{i=1}^{k}(2k_{i}-n_{i}),\geqslant d))_{\mathfrak{q}}$, where $d\geqslant\min\{d_{1},d_{2},\ldots,d_{k}\}$.
Proof.
By Corollary 3, $\mathscr{C}_{i}$ is dual-preserving $f_{i}(x)$-polycyclic code over the chain ring $\mathscr{R}_{i}$ for all $i$. Suppose that $\big{(}g_{1}(x),g_{2}(x),\ldots,g_{k}(x)\big{)}$ and $\big{(}h_{1}(x),h_{2}(x)\ldots,h_{k}(x)\big{)}$ are the images of $g(x)$ and $h(x)$ under the mapping $\eta$ defined in Eq. (7), respectively. Since $g(x)=\oplus_{i=1}^{k}g_{i}(x)e_{i}$ and $h(x)=\oplus_{i=1}^{k}h_{i}(x)e_{i}$, we have $f_{i}(x)=g_{i}(x)h_{i}(x)$ for all $i$. Moreover, since $f_{0}$ is invertible, the constant term of the polynomial $f_{i}(x)$ is invertible for all $i$. By Corollary 4, there exists an $((n_{i},2k_{i}-n_{i},\geqslant d_{i}))$ stabilizer code for all $i$. It is well-know that the direct sum of two quantum-error-correcting codes $((n,k,d))_{\mathfrak{q}}$ and $((n^{\prime},k^{\prime},d^{\prime}))_{\mathfrak{q}}$ is an $((n+n^{\prime},k+k^{\prime},d^{\prime\prime}))_{\mathfrak{q}}$ code, where $d^{\prime\prime}=\min\{d,d^{\prime}\}$, see [8] Section 4. Therefore, there exists a $((\sum_{i=1}^{k}n_{i},\sum_{i=1}^{k}(2k_{i}-n_{i}),\geqslant d))_{\mathfrak{q}}$ stabilizer code, where $d\geqslant\min\{d_{1},d_{2},\ldots,d_{k}\}$. ∎
8 Conclusion

### v3

@@ 7 Annihilator CSS construction
The Euclidean CSS construction is applied to classical linear codes that preserve Euclidean duals to get quantum codes, see for example [34]. In this section, we aim to introduce the annihilator CSS construction for polycyclic codes based on the annihilator dual.
Let $R$ be a finite commutative Frobenius ring with $\mathfrak{q}$ elements. Let $f(x)=x^{n}-\sum_{r=0}^{n-1}f_{r}x^{r}\in R[x]$ and $C$ be an $f(x)$-polycyclic code. According to Section 6, the annihilator product $\langle g(x),h(x)\rangle_{f}$ is defined as $r(0)$, where $r(x)$ is the reminder of $gh(x)$ divided by $f$. The annihilator dual with respect to this product is denoted by $C^{\circ}.$ The Gram matrix $A$ associated with the polynomial $f$ is a $n\times n$ matrix with entries $a_{i,j}=\langle x^{i-1},x^{j-1}\rangle_{f}$. For example, if $f(x)=x^{5}-f_{4}x^{4}-f_{3}x^{3}-f_{2}x^{2}-f_{1}x-f_{0}$, then
$A=\begin{pmatrix}1&0&0&0&0\\[4.0pt] 0&0&0&0&f_{0}\\[4.0pt] 0&0&0&f_{0}&f_{0}f_{4}\\[4.0pt] 0&0&f_{0}&f_{0}f_{4}&f_{0}(f_{4}^{2}+f_{3})\\[4.0pt] 0&f_{0}&f_{0}f_{4}&f_{0}(f_{4}^{2}+f_{3})&f_{0}\left(f_{4}^{3}+2f_{4}f_{3}+f_{2}\right)\end{pmatrix}.$ (21)
The Gram matrix $A$ is symmetric, and if $f_{0}$ is invertible, then $A$ is also invertible. The following lemma provides a recursive formula for calculating the matrix $A$.
Lemma 3.
Let $f(x)=x^{n}-\sum_{r=0}^{n-1}f_{r}x^{r}\in R[x]$ and $A=(a_{i,j})$ be an $n\times n$ Gram matrix associated with $f$. Then $a_{i,j}\;=\;c_{\,i+j-2}$, where
$c_{k}=\begin{cases}1,&k=0,\\ 0,&0<k<n,\\ \sum_{r=0}^{n-1}f_{r}\,c_{k-n+r},&k\geq n.\end{cases}$
Proof.
To compute each $a_{i,j}$, it suffices to find the constant term of the associated polynomial
$x^{\,i+j-2}\bigr{\rvert}_{x^{n}=f_{n-1}x^{n-1}+\cdots+f_{1}x+f_{0}}.$
Let us denote the constant term of $x^{k}|_{x^{n}}$ by $c_{k}$. Therefore, $A$ is the following Hankel matrix [^1] A matrix is called Hankel if each ascending skew-diagonal from left to right is constant.
$A=\begin{pmatrix}c_{0}&c_{1}&c_{2}&\cdots&c_{n-2}&c_{n-1}\\ c_{1}&c_{2}&c_{3}&\ldots&c_{n-1}&c_{n}\\ c_{2}&c_{3}&c_{4}&\ldots&c_{n}&c_{n+1}\\ \vdots&\vdots&\vdots&\ddots&\vdots&\vdots\\ c_{n-2}&c_{n-1}&c_{n}&\ldots&c_{2n-4}&c_{2n-3}\\ c_{n-1}&c_{n}&c_{n+1}&\ldots&c_{2n-3}&c_{2n-2}\end{pmatrix}.$
It is clear that $c_{k}=1$ if $k=0$, and $c_{k}=0$ if $0<k<n$. For $k\geqslant n$, we have $x^{k}=x^{k-n}x^{n}=x^{k-n}\sum_{r=0}^{n-1}f_{r}\,x^{r}=\sum_{r=0}^{n-1}f_{r}\,x^{(k-n)+r}$. Therefore,
$c_{k}=\left(\text{constant term of }x^{k}\bigr{\rvert}_{x^{n}}\right)=\sum_{r=0}^{n-1}f_{r}\left(\left.\text{constant term of }x^{(k-n)+r}\right|_{x^{n}}\right)=\sum_{r=0}^{n-1}f_{r}\,c_{\,k-n+r}.$
Lemma 4.
If $R$ is a commutative ring and $f_{0}$ is invertible, then $C^{\circ}=(CA)^{\perp}=C^{\perp}A^{-1}$, where $CA=\{cA\,|\,c\in C\}$.
Proof.
[…]
Remark 6.
Suppose that $f(x)=x^{n}-f_{0}\in R[x]$ with $f_{0}$ invertible. Then $A$ and its inverse are the monomial matrices
$A=\begin{pmatrix}1&0&\ldots&0&0\\ 0&0&\ldots&0&f_{0}\\ 0&0&\ldots&f_{0}&0\\ \vdots&\vdots&\iddots&\vdots&\vdots\\ 0&f_{0}&\ldots&0&0\end{pmatrix},\qquad A^{-1}=\begin{pmatrix}1&0&\ldots&0&0\\ 0&0&\ldots&0&f_{0}^{-1}\\ 0&0&\ldots&f_{0}^{-1}&0\\ \vdots&\vdots&\iddots&\vdots&\vdots\\ 0&f_{0}^{-1}&\ldots&0&0\end{pmatrix}.$
In this case, since $C^{\circ}=C^{\perp}A^{-1}$, the dual codes $C^{\circ}$ and $C^{\perp}$ are monomially equivalent, which ensures they have the same minimum Hamming distance. However, in general, $A$ may not preserve the minimum Hamming distance because it may not preserve the Hamming weight.
Below is a well-known theorem for constructing quantum codes from classical linear codes that preserve the Euclidean dual. Note that $[[n,k,d]]_{\mathfrak{q}}$ will denote a $R$-ary quantum code of length $n$, with $\mathfrak{q}^{k}$ codewords, and the minimum distance $d$.
Theorem 13 (Euclidean CSS construction [34]).
Let $C_{1}$ and $C_{2}$ be two linear codes over the finite commutative Frobenius ring $R$ with parameters $[n,k_{1},d_{1}]_{R}$ and $[n,k_{2},d_{2}]_{R}$, respectively. If $C_{2}^{\perp}\subseteq C_{1}$, then there exists a quantum stabilizer code with parameters $[[n,k_{1}+k_{2}-n,d]]_{\mathfrak{q}}$, where $d=\min\{\operatorname{wt}(c)\mid c\in(C_{1}\setminus C_{2}^{\perp})\cup(C_{2}\setminus C_{1}^{\perp})\}$ is the minimum Hamming distance.
Now, we are ready to introduce the CSS construction for polycyclic codes that preserve annihilator duals.
@@
Suppose that $R$ is a Frobenius ring and $f_{0}\in R$ is invertible. Let $f(x)=g_{1}(x)h_{1}(x)=g_{2}(x)h_{2}(x)$, and let $C_{1}=\langle g_{1}(x)\rangle$ and $C_{2}=\langle g_{2}(x)\rangle$ be two $f(x)$-polycyclic codes over $R$ with parameters $[n,k_{1},d_{1}]_{R}$ and $[n,k_{2},d_{2}]_{R}$, respectively. If $C^{\circ}_{2}\subseteq C_{1}$ then there exists a quantum stabilizer code with parameters $[[n,k_{1}+k_{2}-n,d]]_{\mathfrak{q}}$, where $d=\min\{\operatorname{wt}(c)\mid c\in(C_{1}\setminus C_{2}^{\circ})\cup(C_{2}A\setminus C_{1}^{\circ}A)\}$ is the minimum Hamming distance.
@@
Since the $n\times n$ Gram matrix $A$ is invertible over $R$, then $xA=0$ leads to $x=0$ for any $x\in R^{n}$ (similar to Corollary 2.4 in [27]). So the transformation $x\mapsto xA$ is a bijective over $R^{n}$, which implies that the dimension of the code $C_{2}A$ equals to $k_{2}$. So, if we replace $C_{2}$ by $C_{2}A$ in the Theorem above, then there exists a quantum stabilizer code with parameters $[[n,k_{1}+k_{2}-n,d]]_{\mathfrak{q}}$, where $d=\min\{\operatorname{wt}(c)\mid c\in(C_{1}\setminus(C_{2}A)^{\perp})\cup(C_{2}A\setminus C_{1}^{\perp})\}$. Applying Lemma 4, we obtain $d=\min\{\operatorname{wt}(c)\mid c\in(C_{1}\setminus C_{2}^{\circ})\cup(C_{2}A\setminus C_{1}^{\circ}A)\}$. ∎
@@
Suppose that $R$ is a finite Frobenius ring and $f_{0}\in R$ is invertible. Let $f(x)=g(x)h(x)$, and let $C=\langle g(x)\rangle$ be an $f(x)$-polycyclic code over $R$ with parameters $[n,k,d_{1}]_{R}$. If $C^{\circ}\subseteq C$, then there exists an $[[n,2k-n,d]]_{\mathfrak{q}}$ stabilizer code, where $d=\min\{\operatorname{wt}(c)\mid c\in(C\setminus C^{\circ})\cup(CA\setminus C^{\circ}A)\}$.
Example 7.1.
Consider the serial ring $\mathscr{R}=\mathbb{F}_{3}[x_{1},x_{2}]/\langle x_{1}^{2}-x_{1},x_{2}^{2}-x_{2}\rangle$, which is isomorphic to $\mathbb{F}_{3}+u\mathbb{F}_{3}+v\mathbb{F}_{3}+uv\mathbb{F}_{3}$ with $u^{2}=u$, $v^{2}=v$, and $uv=vu$. According to Proposition 2, $\mathscr{R}$ is a Frobenius ring. Let $f(x)=x^{5}+2x^{3}+x^{2}+2=(x+1)^{4}(x+2)\in\mathscr{R}[x]$. The $f(x)$-polycyclic code $\mathscr{C}$ generated by $g(x)=x+1$ is a free $\mathscr{R}$-module of dimension $k=\deg(f)-\deg(g)$ with the generating set $\{g(x),xg(x),\ldots,x^{k-1}g(x)\}$. Consider the Gray map $\varphi:\mathscr{R}\rightarrow\mathbb{F}_{3}^{4}$ defined as $\varphi(a+ub+vc+uvd)=(a,a+b,a+c,a+b+c+d)$ in [5]. Using equation (21), one can compute the Gram matrix $A$. Aplying Theorem 12, $\mathscr{C}$ is annihilator dual-containing. Using Magma [11], we compute that $d=2$. Therefore, we obtain a $[[5,3,2]]_{3^{4}}$ quantum stabilizer code.
8 Conclusion and future work